OpenStudy (anonymous):

determine the radius of convergence of given power series summation notation n=0 nx^n/4^n

OpenStudy (anonymous):

\[ a_n=\frac{ n x^n}{4^n}\\ \frac { a_{n+1}}{a_n}=\frac {( n+1) x^{n+1}}{4^{n+1}} \frac {4^n} {n x^n}=\frac 1 4 \frac {n+1}{n}x\to \frac x 4 \] radius of convergence r=4

OpenStudy (anonymous):

You have to take the limit of the absolute value of the ratios in my previous post and the limit will be \[ \frac {|x|} 4 \]

OpenStudy (anonymous):

thank u very much. your answer was correct. i have another one if u want to try determin the taylor series about point x subscript o= 5 for the function x^2